### v1

@@ 4 Quantifying Privacy Leakage with PANORAMIA
The first step to quantifying privacy leakage is to formalize an auditing game. PANORAMIA starts with $x^{\textrm{in}}\in{\mathcal{X}}^{m}$ training points from ${\mathcal{D}}$ (e.g., the data distribution of one participant in an FL setting), as well as $x^{\textrm{gen}}\in{\mathcal{X}}^{m}$ generated points from the generator distribution ${\mathcal{G}}$ ($x^{\textrm{gen}}\sim{\mathcal{G}}$). We construct a sequence auditing samples $x\in{\mathcal{X}}^{m}$ as follows:
Definition 2 (Auditing game).
[…]
This means that $s$ is sampled independently to choose either the real ($s_{i}=1$) or generated ($s_{1}=0$) data point at each index $i$. This creates a sequence of $m$ examples that PANORAMIA will try to tell apart (i.e., guess $s$). The level of success achievable in this game quantifies the privacy leakage of the target model $f$. More precisely, we follow an analysis inspired by that of Steinke et al. (2023), but require several key technical changes to support auditing with no retraining using generated non-member data. To realize this, we first formalize PANORAMIA’s audit procedure as a hypothesis test on our auditing game, for which we construct a statistical test (§4.1). Then, we show how to use this hypothesis test to quantify privacy leakage as a lower confidence interval, and interpret the semantics of our privacy leakage measurement (§4.2).
4.1 Formalizing the Audit as a Hypothesis Test
@@
The smaller $c$, the better generator, a ${\mathcal{G}}$ assigns a probability to real data that cannot be much smaller than that of the real data distribution. We make two important remarks.
@@
Using Definition 3, we can formulate the hypothesis test on the auditing game of Definition 2 that underpins our approach:
@@
Note that, in a small abuse of language, we will often refer to $f$ as $\epsilon$-DP to say that $f$ is the output of an $\epsilon$-DP mechanism.
To construct a statistical test allowing us to reject ${\mathcal{H}}$ based on evidence, we define two key mechanisms (corresponding to PANORAMIA’s auditing scheme). First, the (potentially randomized) baseline guessing mechanism $B(s,x):\{0,1\}^{m}\times{\mathcal{X}}^{m}\rightarrow{\mathbb{R}}_{+}^{m}$, which outputs a (non-negative) score for the membership of each datapoint $x_{i}$, based on this datapoint only. That is, $B(s,x)=\{b(x_{1}),b(x_{2}),\ldots,b(x_{m})\}$.
Second, we define $A(s,x,f):\{0,1\}^{m}\times{\mathcal{X}}^{m}\times{\mathcal{F}}\rightarrow{\mathbb{R}}_{+}^{m}$, which outputs a (potentially randomized) non-negative score for the membership of each datapoint, with the guess for index $i$ depending on $x_{\leq i}$ and target model $f$. Note that if the target model $f$ is DP, then $A$ is DP w.r.t. inclusion in the auditing game $s$, outside of what is revealed by $x$. We are now ready to construct a hypothesis test for ${\mathcal{H}}$. First, we construct tests for each part of the hypothesis separately.
@@
${\mathbb{P}}_{S,X,T^{b}}\Big{[}\sum_{i=1}^{m}T^{b}_{i}\cdot S_{i}\geq v\ |\ T^{b}=t^{b}\Big{]}$
[…]
@@
${\mathbb{P}}_{S,X,T^{a}}\Big{[}\sum_{i=1}^{m}T^{a}_{i}\cdot S_{i}\geq v\ |\ T^{a}=t^{a}\Big{]}\leq$
[…]
@@
${\mathbb{P}}\Big{[}\sum_{i=1}^{m}T^{a}_{i}\cdot S_{i}\geq v^{a},\ \sum_{i=1}^{m}T^{b}_{i}\cdot S_{i}\geq v^{b}\ |\ T^{a}=t^{a},T^{b}=t^{b}\Big{]}$
$\leq\beta^{a}(m,c,\epsilon,v^{a},t^{a})+\beta^{b}(m,c,v^{b},t^{b})$
@@
$\underset{{S^{\prime}\sim\textrm{Bernoulli}(p)^{r}}}{{\mathbb{P}}}\Big{[}\sum_{i=1}^{r}\frac{S^{\prime}_{i}}{r}\geq p+\sqrt{\frac{\log(1/\beta)}{2r}}\Big{]}\leq\beta,$
@@ 4.2 Quantifying Privacy Leakage and Interpretation
In an audit we would like to quantify $\epsilon$, not just reject a given $\epsilon$ claim. We use the hypothesis test from Corollary 1 to compute a confidence interval on $c$ and $\epsilon$. To do this, we first define an ordering between $(c,\epsilon)$ pairs, such that if $(c_{1},\epsilon_{1})\leq(c_{2},\epsilon_{2})$, the event (i.e., set of observations for $T^{a,b},S$) for which we can reject ${\mathcal{H}}(c_{2},\epsilon_{2})$ is included in the event for which we can reject ${\mathcal{H}}(c_{1},\epsilon_{1})$. That is, if we reject ${\mathcal{H}}$ for values $(c_{2},\epsilon_{2})$ based on audit observations, we also reject values $(c_{1},\epsilon_{1})$ based on the same observations.
We define the following lexicographic order to fit this assumption, based on the hypothesis test from Corollary 1:
@@
With this ordering, we have that:
@@
[…]
}$\big{\}}\Big{]}\geq 1-\beta$
@@
Apply Lemma 4.7 from Steinke et al. (2023) with the ordering from Eq. 3 and the test from Corollary 1. ∎
@@
The lower confidence interval for $(c,\epsilon)$ at confidence $1-\beta$ is the largest $(c,\epsilon)$ pair that cannot be rejected using Corollary 1 with false rejection probability at most $\beta$. Hence, for a given confidence level $1-\beta$, PANORAMIA computes $(c_{\textrm{lb}},\tilde{\epsilon})$, the largest value for $(c,\epsilon)$ that it cannot reject. Note that Corollaries 1 and 2 rely on a union bound between two tests, one for $c$ and one for $c+\epsilon$. We can thus consider each test separately (at level $\beta/2$). In practice, we follow previous approaches Maddock et al. (2023); Steinke et al. (2023) and use the whole precision/recall curve. Each level of recall (i.e., threshold on $t^{a,b}$ to predict membership) corresponds to a bound on the precision, which we can compare to the empirical value. Separately for each test, we pick the level of recall yielding the highest lower-bound (Lines 4-7 in the last section of Algorithm 1). We next discuss the semantics of returned values, $(c_{\textrm{lb}},\tilde{\epsilon})$.
@@
Audit semantics. Corollary 2 gives us a lower-bound for $(c,\epsilon)$, based on the ordering from Equation 3. To understand the value $\tilde{\epsilon}$ returned by PANORAMIA, we need to understand what the hypothesis test rejects. Rejecting ${\mathcal{H}}$ means either rejecting the claim about $c$, or the claim about $c+\epsilon$ (which is the reason for the ordering in Equation 3). With Corollary 2, we hence get both a lower-bound $c_{\textrm{lb}}$ on $c$, and $\{c\!+\!\epsilon\}_{\textrm{lb}}$ on $c+\epsilon$. Unfortunately, $\tilde{\epsilon}\triangleq\max\{0,\{c\!+\!\epsilon\}_{\textrm{lb}}-c_{\textrm{lb}}\}$, which is the value PANORAMIA returns, does not imply a lower-bound on $\epsilon$. Instead, we can claim that “PANORAMIA could not reject a claim of $c$-closeness for ${\mathcal{G}}$, and if this claim is tight, then $f$ cannot be the output of a mechanism that is $\tilde{\epsilon}$-DP”.
@@
While this is not as strong a claim as typical lower-bounds on $\epsilon$-DP from prior privacy auditing works, we believe that this measure is useful in practice. Indeed, the $\tilde{\epsilon}$ measured by PANORAMIA is a quantitative privacy measurement, that is accurate (close to a lower-bound on $\epsilon$-DP) when the baseline performs well (and hence $c_{\textrm{lb}}$ is tight). Thus, when the baseline is good, $\tilde{\epsilon}$ can be interpreted as (close to) a lower bound on (pure) DP. In addition, since models on the same dataset and task share the same baseline, which does not depend on the audited model, PANORAMIA’s measurement can be used to directly compare privacy leakage between models. Thus, PANORAMIA opens a new capability, measuring privacy leakage of a trained model $f$ without access or control of the training pipeline, with an meaningfull and practically useful measurement, as we empirically show next.
@@ 5 Evaluation
ResNet101 CIFAR10 20, 50, 100 91.61%, 90.18%, 87.93% ResNet101-E20, ResNet101-E50, ResNet101-E100,
@@
GPT-2 WikiText-2 12, 46, 92 Refer to Appendix B.2 GPT-2_E12, GPT-2_E46, GPT-2_E92
MLP Tabular Classification Adult 10, 100 86%, 82% MLP_E10, MLP_E100
Table 2: Train and Test Metrics for ML Models Audited. *”Model Variants” trained for different number of epochs $E$.
We instantiate PANORAMIA on target models for four tasks from three data modalities. For image classification, we consider the CIFAR10 Krizhevsky et al. (2009), and CelebA Liu et al. (2015) datasets, with varied target models: a four-layers CNN O’Shea & Nash (2015), a ResNet101 He et al. (2015) and a differentially-private ResNet18 He et al. (2015) trained with DP-SGD (Abadi et al., 2016) using Opacus (Yousefpour et al., 2021) at different values of $\epsilon$. We use StyleGAN2 Karras et al. (2020) for ${\mathcal{G}}$. For language models, we fine-tune small GPT-2 Radford et al. (2019) on the WikiText-2 train dataset Merity et al. (2016) (we also incorporate documents from WikiText-103 to obtain a larger dataset). ${\mathcal{G}}$ is again based on small GPT-2, and then fine-tuned on $D_{G}$. We generate samples using top-$p$ sampling Holtzman et al. (2019) and a held-out prompt dataset $D_{G}^{prompt}\subset D_{G}$. Finally, for classification on tabular data, we fit a Multi-Layer Perceptron (MLP) with 4 hidden layers trained on the Adult dataset Becker & Kohavi (1996), on a binary classification task predicting income >\$$>\$50k$. We use the MST algorithm McKenna et al. (2021) for ${\mathcal{G}}$.
Table 2 summarizes the tasks, models and performance, as well as the respective names we use to show results. Due to lack of space, more details on the settings, the models used and implementation details are provided in  Appendix B. Our results are organized as follows. First, we evaluate the strength of our baseline (§5.1), on which the semantics of PANORAMIA’s audit rely. Second, we show what PANORAMIA detects meaningful privacy leakage in our settings, comparable to the lower-bounds provided by the O(1) approach (Steinke et al., 2023) (though under weaker requirements) (§5.2). Finally, we show that PANORAMIA can detect varying amounts of leakage from models with controlled data leakage using model size and DP (§5.3).
@@ 5.1 Baseline Design and Evaluation
Our MIA is a loss-based attack based on an attack model that takes as input a datapoint $x$ as well as the value of the loss of target model $f$ on point $x$. Appendix B details the architectures used for the attack model for each data modality. Recall from §4.2 the importance of having a tight $c_{\textrm{lb}}$ for our measure $\tilde{\epsilon}$ to be close to a lower-bound on $\epsilon$-DP, which also requires a strong baseline. To increase the performance of our baseline $b$, we mimic the role of the target model $f$’s loss in the MIA using a helper model $h$, which adds a loss-based feature to $b$.
This new feature can be viewed as side information about the data distribution. Table 3 shows the $c_{\textrm{lb}}$ value under different designs for $h$. The best performance is consistently when $h$ is trained on synthetic data before being used as a feature to train the $b$. Indeed, such a design reaches a $c_{\textrm{lb}}$ up to $1.36$ larger that without any helper (CIFAR10) and $0.16$ higher than when training on real non-member data without requiring access to real non-member data, a key requirement in PANORAMIA. We adopt this design in all the following experiments. In Figure 2 we show that the baseline has enough training data (vertical dashed line) to reach its best performance. Section C.1, shows similar results on all modalities (except on tabular data, in which more data for $b$ would help a bit), as well as different architectures that we tried. All these evidences confirm the strength of our baseline.
@@
[Uncaptioned image] Figure 2: CIFAR-10 baseline on increasing training size.
5.2 Main Auditing Results
We run PANORAMIA on models with different values of over-fitting (by varying the number of epochs, see the final accuracy on Table 2) for each data modality. More over-fitted models are known to leak more information about their training data due to memorization Yeom et al. (2018); Carlini et al. (2019; 2022a). To show the auditing power of PANORAMIA, we compare it with two strong approaches to lower-bounding privacy loss. First, we use a variation of our approach using real non-member data instead of generated data (called RM;RN for Real Members; Real Non-members). While this basically removes the role of the baseline ($c_{\textrm{lb}}=0$), it requires access to a large sample of non-member data from the same distributions as members (hard requirement) or the possibility of training costly shadow models to create such non-members. Second, we rely on the O(1) audit from Steinke et al. (2023), which is similar to the previous approach but predicts membership based on a loss threshold while also leveraging guesses on non-members in its statistical test. Note that this technique requires control of the training process. The privacy loss measured by these techniques gives a target what we hope PANORAMIA to detect.

### v2

@@ 4 Quantifying Privacy Leakage with PANORAMIA
We formalize a privacy game as follows: construct a sequence of auditing samples, in which $s_{i}$ is sampled independently to choose either the real ($x^{\textrm{in}}\in{\mathcal{X}}^{m}$ training points from ${\mathcal{D}}$ if $s_{i}=1$) or generated ($x^{\textrm{gen}}\in{\mathcal{X}}^{m}$ generated points from ${\mathcal{G}}$ if $s_{1}=0$) data point at each index $i$.
Definition 2 (Privacy game).
$s\sim\textrm{Bernoulli}(\frac{1}{2})^{m},\ \textrm{with}\ s_{i}\in\{0,1\},x_{i}=(1-s_{i})x^{\textrm{gen}}_{i}+s_{i}x^{\textrm{in}}_{i},\ \forall i\in\{1,\ldots,m\}.$
The key difference between Definition 2 and the privacy game of Steinke et al. (2023) lies in how we create the audit set. Instead of sampling non-members independently and removing them from the training set (which requires modifyig the training set and the model), Definition 2 starts from a set of known members (after the fact), and pairs each point with a non-member (generated i.i.d. from the generator distribution). For each pair, we then flip a coin to decide which point in the pair will be shown to the auditor for testing, thereby creating the test task of our privacy measurement.
The level of achievable success in this game quantifies the privacy leakage of the target model $f$. More precisely, we follow an analysis inspired by that of Steinke et al. (2023), but require several key technical changes to support auditing with no retraining using generated non-member data.
4.1 Formalizing the Privacy Measurements as a Hypothesis Test
@@
The smaller $c$, the better the generator, as ${\mathcal{G}}$ assigns a probability to real data that cannot be much smaller than that of the real data distribution. We make three important remarks.
@@
Remark 2: We show in Appendix E that we can relax this definition to hold only with high probability (following the $(\epsilon,\delta)$-DP relaxation). Empirical results under this relaxation are very similar. We also show how to measure $(\epsilon,\delta)$-DP, instead of $\epsilon$-DP on which we focus here.
@@
Using Definition 3, we can formulate the hypothesis test on the privacy game of Definition 2 that underpins our approach:
@@
Note that, in a small abuse of language, we will often refer to $f$ as $\epsilon$-DP to say that $f$ is the output of an $\epsilon$-DP training mechanism. That is, $B(s,x)=\{b(x_{1}),b(x_{2}),\ldots,b(x_{m})\}$.
We define two key mechanisms: $B(s,x):\{0,1\}^{m}\times{\mathcal{X}}^{m}\rightarrow{\mathbb{R}}_{+}^{m}$ outputs a (potentially randomized) non-negative score for the membership of each datapoint $x_{i}$, based on $x_{i}$ only; $A(s,x,f):\{0,1\}^{m}\times{\mathcal{X}}^{m}\times{\mathcal{F}}\rightarrow{\mathbb{R}}_{+}^{m}$ outputs a (potentially randomized) non-negative score for the membership of each datapoint, with the guess for index $i$ depending on $x_{\leq i}$ and target model $f$. We can construct a statistical test for each part of the hypothesis separately.
@@
${\mathbb{P}}_{S,X,T^{b}}\Big{[}\sum_{i=1}^{m}T^{b}_{i}\cdot S_{i}\geq v\ |\ T^{b}=t^{b}\Big{]}\leq\underset{{S^{\prime}\sim\textrm{Bernoulli}(\frac{e^{c}}{1+e^{c}})^{m}}}{{\mathbb{P}}}\Big{[}\sum_{i=1}^{m}t^{b}_{i}\cdot S^{\prime}_{i}\geq v\Big{]}\triangleq\beta^{b}(m,c,v,t^{b})$
@@
${\mathbb{P}}_{S,X,T^{a}}\Big{[}\sum_{i=1}^{m}T^{a}_{i}\cdot S_{i}\geq v\ |\ T^{a}=t^{a}\Big{]}\leq\underset{{S^{\prime}\sim\textrm{Bernoulli}(\frac{e^{c+\epsilon}}{1+e^{c+\epsilon}})^{m}}}{{\mathbb{P}}}\Big{[}\sum_{i=1}^{m}t^{a}_{i}\cdot S^{\prime}_{i}\geq v\Big{]}\triangleq\beta^{a}(m,c,\epsilon,v,t^{a})$
@@
${\mathbb{P}}\Big{[}\sum_{i=1}^{m}T^{a}_{i}\cdot S_{i}\geq v^{a},\ \sum_{i=1}^{m}T^{b}_{i}\cdot S_{i}\geq v^{b}\ |\ T^{a}=t^{a},T^{b}=t^{b}\Big{]}\leq\beta^{a}(m,c,\epsilon,v^{a},t^{a})+\beta^{b}(m,c,v^{b},t^{b})$
@@
$\underset{{S^{\prime}\sim\textrm{Bernoulli}(p)^{r}}}{{\mathbb{P}}}\Big{[}\sum_{i=1}^{r}\frac{S^{\prime}_{i}}{r}\geq p+\sqrt{\frac{\log(1/\beta)}{2r}}\Big{]}\leq\beta,$ (3)
@@ 4.2 Quantifying Privacy Leakage and Interpretation
In an privacy measurement, we would like to quantify $\epsilon$, not just reject a given $\epsilon$ claim. We use the hypothesis test from Corollary 1 to compute a confidence interval on $c$ and $\epsilon$. To do this, we first define an ordering between $(c,\epsilon)$ pairs, such that if $(c_{1},\epsilon_{1})\leq(c_{2},\epsilon_{2})$, the event (i.e., set of observations for $T^{a,b},S$) for which we can reject ${\mathcal{H}}(c_{2},\epsilon_{2})$ is included in the event for which we can reject ${\mathcal{H}}(c_{1},\epsilon_{1})$. That is, if we reject ${\mathcal{H}}$ for values $(c_{2},\epsilon_{2})$ based on audit observations, we also reject values $(c_{1},\epsilon_{1})$ based on the same observations.
We define the following order to fit this assumption, based on the hypothesis test from Corollary 1:
@@
This lexicographic order over hypotheses ensures that, when using the joint hypothesis test from Corollary 1 to construct confidence intervals, we always reject hypotheses with a low value of $c$ incompatible with the baseline performance. Formally, this yields the following confidence intervals:
@@
For all $\beta\in]0,1]$, $m$, and observed $t^{a},t^{b}$, call $v_{\textrm{ub}}^{a}(c,\epsilon)\triangleq\sup\ \{v:\beta^{a}(m,c,\epsilon,v,t^{a})\leq\frac{\beta}{2}\}$ and $v_{\textrm{ub}}^{b}(c)\triangleq\sup\ \{v:\beta^{b}(m,c,v,t^{b})\leq\frac{\beta}{2}\}$. Then:
${\mathbb{P}}\Big{[}(c,\epsilon)\geq\sup\big{\{}(c^{\prime},\epsilon^{\prime}):t^{a}\cdot s\leq v_{\textrm{ub}}^{a}(c^{\prime},\epsilon^{\prime})\ \textrm{and}\ t^{b}\cdot s\leq v_{\textrm{ub}}^{b}(c^{\prime})\big{\}}\Big{]}\geq 1-\beta$
@@
Apply Lemma 4.7 from Steinke et al. (2023) with the ordering from Equation 4 and the test from Corollary 1. ∎
@@
The lower confidence interval for $(c,\epsilon)$ at confidence $1-\beta$ is the largest $(c,\epsilon)$ pair that cannot be rejected using Corollary 1 with false rejection probability at most $\beta$. Hence, for a given confidence level $1-\beta$, PANORAMIA computes $(c_{\textrm{lb}},\tilde{\epsilon})$, the largest value for $(c,\epsilon)$ that it cannot reject. Note that Corollaries 1 and 2 rely on a union bound between two tests, one for $c$ and one for $c+\epsilon$. We can thus consider each test separately (at level $\beta/2$). In practice, we use the whole precision/recall curve to achieve tighter bounds: each level of recall (i.e., threshold on $t^{a,b}$ to predict membership) corresponds to a bound on the precision, which we can compare to the empirical value. Separately for each test, we pick the level of recall yielding the highest lower-bound (Lines 4-7 in the last section of Algorithm 1). For all bounds to hold together, we apply a union bound over all tests at each level of recall. Specially, if we have $m$ test datapoints, we have at most $m$ different values of recall, and the union bound corresponds to $\beta\leftarrow\beta/m$ (notice in Equation 3 that the bounds depend on $\sqrt{\log(1/\beta)}$, so this is not too costly). We next discuss the semantics of returned measurement values $(c_{\textrm{lb}},\tilde{\epsilon})$.
@@
Measurement semantics. Corollary 2 gives us a lower-bound for $(c,\epsilon)$, based on the ordering from Equation 4. To understand the value $\tilde{\epsilon}$ returned by PANORAMIA, we need to understand what the hypothesis test rejects. Rejecting ${\mathcal{H}}$ means either rejecting the claim about $c$, or the claim about $c+\epsilon$ (which is the reason for the ordering in Equation 4). With Corollary 2, we hence get both a lower-bound $c_{\textrm{lb}}$ on $c$, and $\{c\!+\!\epsilon\}_{\textrm{lb}}$ on $c+\epsilon$. Unfortunately, $\tilde{\epsilon}\triangleq\max\{0,\{c\!+\!\epsilon\}_{\textrm{lb}}-c_{\textrm{lb}}\}$, which is the value PANORAMIA returns, does not imply a lower-bound on $\epsilon$. Instead, we can claim that “PANORAMIA could not reject a claim of $c$-closeness for ${\mathcal{G}}$, and if this claim is tight, then $f$ cannot be the output of a mechanism that is $\tilde{\epsilon}$-DP”.
@@
While this is not as strong a claim as typical lower-bounds on $\epsilon$-DP from prior privacy auditing works, we believe that this measure is useful in practice. Indeed, the $\tilde{\epsilon}$ measured by PANORAMIA is a quantitative privacy measurement, that is accurate (close to a lower-bound on $\epsilon$-DP) when the baseline performs well (and hence $c_{\textrm{lb}}$ is tight). Thus, when the baseline is good, $\tilde{\epsilon}$ can be interpreted as (close to) a lower bound on (pure) DP. PANORAMIA opens a new capability, measuring privacy leakage of a trained model $f$ without access or control of the training pipeline, with a meaningful and practically useful measurement, as we empirically show next.
@@ 5 Evaluation
ResNet101 CIFAR10 20, 50, 100 91.61%, 90.18%, 87.93% ResNet101_E20, ResNet101_E50, ResNet101_E100,
WideResNet-28-10 CIFAR10 150, 300 95.67%, 94.23% WRN-28-10_E150, WRN-28-10_E300
ViT-small (pretrained) CIFAR10 35 96.38% ViT_E35
@@
GPT-2 WikiText 37, 75, 150 Refer to Appendix C.2 GPT-2_E37, GPT-2_E75, GPT-2_E150
Table 1: Train and Test Metrics for ML Models Audited. *"Model Variants" trained for different numbers of epochs $E$. For ViT-small, we use a model pre-trained on imagenet and tune it on CIFAR10.
We instantiate PANORAMIA on target models for four tasks from three data modalities [^1] Code available here: https://github.com/ubc-systopia/panoramia-privacy-measurement.. For image classification, we consider the CIFAR10 Krizhevsky (2009), and CelebA Liu et al. (2015) datasets, with varied target models: a four-layers CNN O’Shea & Nash (2015), a ResNet101 He et al. (2015), a Wide ResNet-28-10  Zagoruyko & Komodakis (2017b), a Vision Transformer (ViT)-small  Dosovitskiy et al. (2021), and a DP ResNet18 He et al. (2015) trained with DP-SGD (Abadi et al., 2016) using Opacus (Yousefpour et al., 2021) at different values of $\epsilon$. We use StyleGAN2 Karras et al. (2020) for ${\mathcal{G}}$. For language models, we fine-tune small GPT-2 Radford et al. (2019) on the WikiText train dataset Merity et al. (2016) (we take a subset of the documents in WikiText-103). ${\mathcal{G}}$ is again based on small GPT-2, and then fine-tuned on $D_{G}$. We generate samples using top-$k$ sampling Holtzman et al. (2019) and a held-out prompt dataset $D_{G}^{\textnormal{prompt}}\subset D_{G}$. Finally, we conduct experiments on classification on tabular data. However for this data modality, PANORAMIA did not detect any meaningful leakage. More precisely, we have observed significant variance in the audit values returned by PANORAMIA which make the results inconclusive. Nonetheless, we present the results obtained for this task and discuss this issue further in Section D.6. Table 1 summarizes the tasks, models, and performance. More details on the settings, the models used, and implementation details are provided in Appendix C.
@@ 5.1 Baseline Design and Evaluation
Our MIA is a loss-based attack, which uses an ML model taking as input both a datapoint $x$ and the value of the loss of target model $f$ on point $x$. Appendix C details the architectures used for the attack model for each data modality.
@@
Table 2: Baseline evaluation with different helper model scenarios.
Figure 2: CIFAR-10 baseline on increasing training size.
Recall from §4.2 the importance of having a tight $c_{\textrm{lb}}$ for our measure $\tilde{\epsilon}$ to be close to a lower-bound on $\epsilon$-DP. To increase the performance of our baseline $b$, we mimic the role of the target model $f$’s loss in the MIA using a helper model $h$ trained on synthetic data, which adds a loss-based feature to $b$.
This new feature can be viewed as side information about the data distribution. Table 2 shows the $c_{\textrm{lb}}$ value for different designs for $h$. The best performance is consistently when $h$ is trained on synthetic data before being used as a feature to train the $b$. Indeed, such a design reaches a $c_{\textrm{lb}}$ up to $1.19$ larger that without any helper (CIFAR10) and $0.23$ higher than when training on real non-member data, without requiring access to real non-member data, a key requirement in PANORAMIA. We adopt this design in all the following experiments. In Figure 2 we show that the baseline has enough training data (vertical dashed line) to reach its best performance. Section D.1, shows similar results on GPT-2 as well as different architectures that we tried for the helper model. All these pieces of evidence confirm the strength of our baseline.
5.2 Main Privacy Measurement Results
We run PANORAMIA on models with different degrees of over-fitting by varying the number of epochs, (see the final accuracy on Table 1) for each data modality. More over-fitted models are known to leak more information about their training data due to memorization (Yeom et al., 2018; Carlini et al., 2019, 2022a). To show the privacy loss measurement power of PANORAMIA, we compare it with two strong approaches to lower-bounding privacy loss.
First, we compare PANORAMIA with the O(1) audit (Steinke et al., 2023) (in input space, without canaries in a black box setting), which uses a loss threshold MIA to detect both members and non-members. Second, we use a hybrid between O(1) and PANORAMIA, in which we use real non-member data instead of generated data with our ML-based MIA (called RM;RN for Real Members; Real Non-members). In both cases $c_{\textrm{lb}}=0$, but it requires control over training data inclusion and a large set of non-member data. The privacy loss measured by these techniques gives a target that we hope PANORAMIA to detect.
LABEL:fig:eval:comparison_precision shows the precision of $b$ and PANORAMIA at different levels of recall, and LABEL:fig:eval:comparison_privacy the corresponding value of $\{c+\epsilon\}_{\textnormal{lb}}$ (or $c_{\textrm{lb}}$ for $b$). Dashed lines show the maximum value of $\{c+\epsilon\}_{\textnormal{lb}}$ (or $c_{\textrm{lb}}$) achieved (Fig. LABEL:fig:eval:comparison_privacy) (returned by PANORAMIA), and the precision implying these values at different recalls (Fig. LABEL:fig:eval:comparison_precision). Table 3 summarizes those $\{c+\epsilon\}_{\textnormal{lb}}$ (or $c_{\textrm{lb}}$) values, as well as the $\epsilon$ measured by existing approaches. We make two key observations.